OpenStudy (anonymous):

If a sample of a substance contains 9.03 x 1023 representative particles, how many moles does this represent? 0.5 mole 1 mole 1.5 mole 9 mole

OpenStudy (xishem):

Use a conversion factor to convert from number of particles to moles:\[9.03 \times {10^{23}}\ particles*(\frac{1mol}{6.022 \times {10^{23}}\ particles})=1.5mol\]